OpenStudy (anonymous):

find the Pct to the nearest thousandth....games won-15;games played-25

OpenStudy (paki):

pct....?

OpenStudy (paki):

what is this

OpenStudy (anonymous):

I'm assuming you mean percent? And then the percentage of games won? \[\frac{15}{25}*100\]\[\frac{3}{5}*100\]\[\frac{300}{5}\] = 60%